# Compact stellar systems hosting an intermediate mass black hole: magnetohydrodynamic study of inflow-outflow dynamics

### v1

@@ Abstract
Context. Intermediate-mass black holes (IMBHs) have remained elusive in observations, with only a handful of tentative detections to date. One promising location for the existence of IMBHs is in dense stellar clusters such as IRS 13E in the Galactic Center. Such systems are thought to be fed by stellar winds from massive Wolf–Rayet (WR) stars. Understanding the accretion dynamics in wind-fed IMBHs in such clusters is important for predicting observational signatures. These systems are, however, relatively unexplored theoretically. Understanding the interplay between IMBHs and the surrounding stars—especially through the high-velocity, dense stellar winds of WR stars—is therefore essential for clarifying IMBH dynamics within such environments and for explaining why these objects continue to evade unambiguous detection.
@@
Methods. We conduct three-dimensional magnetohydrodynamic and hydrodynamic simulations in which each WR star is represented as a source term in mass, momentum, energy, and magnetic field. A metallicity-dependent cooling function accounts for radiative energy losses. By varying the cluster’s geometry, magnetization, and chemical abundance, we characterize the resulting flow structures, accretion rates, and X-ray luminosity corresponding to the cluster as a whole as well as the IMBH.
Results. In all configurations, the accretion rate onto the IMBH is up to five orders of magnitude lower than the total mass-loss rate from the cluster’s WR stars. High-velocity wind–wind collisions generate turbulent, shock-heated outflows that expel most of the injected gas. While enhanced cooling in high-metallicity runs fosters dense clump formation, these clumps typically do not reach the black hole. The integrated X-ray emission is dominated by colliding stellar winds, rendering the IMBH’s radiative signature elusive. Intermittent, quasi-periodic variations in inflow rates are driven by close stellar passages, leading to short-lived accretion enhancements or flares that nonetheless remain difficult to detect against the dominant wind emission.
Conclusions. Despite continuous mass injection from dense stellar winds, these compact systems exhibit outflow-dominated flows and weak, though strongly variable black-hole accretion, naturally explaining the low detectability of IMBHs in such settings. The observed quasiperiodic or flaring accretion episodes are overwhelmed by the luminous shock-heated winds, making unambiguous observational identification of the IMBH challenging with current X-ray instruments. Nevertheless, our results provide a framework for interpreting current data and developing future observational strategies aimed at unveiling IMBHs in dense stellar systems.
@@ 1 Introduction
Intermediate-mass black holes (IMBHs) with masses in the range $M_{\bullet}\sim 10^{2}-10^{5}$ solar masses ($M_{\odot}$) have largely eluded detection, leaving a noticeable gap in the mass distribution of black holes (see Greene et al., 2020, for a review). At the same time, comprehending their formation, evolution, and occurrence across cosmic history is crucial for understanding both black-hole growth and galaxy evolution. So far there have been a few confirmed detections towards the upper IMBH mass range in lower-mass galaxies (e.g. NGC 4395 with $M_{\bullet}\sim 10^{4}-10^{5}\,M_{\odot}$; Filippenko & Ho, 2003; Peterson et al., 2005; Pandey et al., 2024). Toward the lower mass range, there was a gravitational wave detection of the formation of a $\sim 140\,M_{\odot}$ black hole (GW190521; Abbott et al., 2020), which falls in the pair-instability mass gap (Woosley & Heger, 2021).
There are several ways that an IMBH could form depending on the cosmic time at formation and the surrounding environment, both of which influence its subsequent growth. Three formation channels can be distinguished: (i) primordial/cosmological origin due to the collapse of population III stars (Madau & Rees, 2001) or a direct gas cloud collapse (Loeb & Rasio, 1994; Eisenstein & Loeb, 1995; Begelman et al., 2006; Latif et al., 2013); (ii) consecutive mergers of stellar-mass black holes in dense stellar systems, such as globular clusters (Miller & Hamilton, 2002; Gültekin et al., 2004; Fujii et al., 2024); and (iii) runaway collisions and mergers of massive stars in dense stellar clusters that subsequently collapse into an IMBH (Portegies Zwart & McMillan, 2002). Despite the uncertainties there are several theoretical models that support (iii), in particular for globular clusters (Miller & Hamilton, 2002; Portegies Zwart & McMillan, 2002; Gültekin et al., 2004; Freitag et al., 2006). Once a sufficiently massive IMBH seed forms ($>50\,M_{\odot}$; Miller & Hamilton, 2002), it is expected to be retained within the cluster and undergo only a low-amplitude Brownian-like motion (Chatterjee et al., 2002a, b).
@@
Even granting those assumptions, however, both the occupancy fraction and mass ratio for clusters are uncertain (see e.g. Šubr et al., 2019) due to uncertainties in the initial cluster concentration (Giersz et al., 2015), which influences whether the hypothetical IMBH formation proceeds rapidly or more gradually. Nevertheless, there are several pieces of observational evidence to support the idea that at least some massive stellar clusters host an IMBH. For instance, the observed off-nuclear X-ray sources and transients in galactic halos with peak X-ray luminosities of $\sim 10^{43}\,{\rm erg\,s^{-1}}$ are consistent with the emission of IMBHs (with masses on the order of $\sim 10^{4}\,M_{\odot}-10^{5}\,M_{\odot}$) powered by the tidal disruption of stars, which supports their association with lower-mass stellar systems like stellar clusters or dwarf galaxies (Lin et al., 2018; Jin et al., 2025; Zhang et al., 2025). There is strong evidence that at least two globular clusters in the Milky Way host IMBHs based on dynamical grounds: $\omega$ Centauri (Häberle et al., 2024) and 47 Tucanae (NGC 104, Kızıltan et al. 2017), with IMBH masses of $\gtrsim 8200\,M_{\odot}$ and $2300_{-850}^{+1500}\,M_{\odot}$, respectively, though several studies point out potential caveats and systematic effects that can influence the central mass constraints (Zocchi et al., 2017, 2019; Bañares-Hernández et al., 2025).
@@
Galactic nuclei, like the Galactic Center in our own galaxy, are especially promising regions for detecting IMBHs. Due to the short dynamical relaxation timescale, as many as $\gtrsim 50(M_{\bullet}/150\,M_{\odot})^{1/2}$ IMBHs could be concentrated towards the supermassive black hole (SMBH, Madau & Rees 2001). These black holes would have been formed either primordially or in globular clusters before undergoing a rapid inspiral caused by dynamical friction. The infall rate is expected to decrease with cosmic time as $\propto t^{-1/2}$ and the current rate is estimated to be $\sim 2$ IMBHs per gigayear per galaxy (Madau & Rees, 2001).
@@
In addition, IMBHs up to $\lesssim 10^{4}\,M_{\odot}$ can also form in situ within the dense nuclear star clusters (NSCs) surrounding SMBHs (Neumayer et al., 2020) via black hole–star collisions (Rose et al., 2022; Haas et al., 2025) and black hole-black hole mergers (Fragione et al., 2022). Such IMBHs may be detected indirectly when they perturb the inner accretion flow, causing quasiperiodic accretion and outflow phenomena (Suková et al., 2021; Pasham et al., 2024), through spin-orbit coupling-induced jet precession (Ressler et al., 2024), or occasionally as X-ray sources when they disrupt a star (Lin et al., 2018; Jin et al., 2025) or encounter a denser molecular cloud (Seepaul et al., 2022; Hu et al., 2024). Furthermore, recent observations suggest that some high-velocity compact clouds (HVCCs) in the Galactic center, such as CO–0.40–0.22, may be influenced by gravitational interactions with inactive IMBHs, highlighting their potential role in the dynamics of molecular clouds (Takekawa et al., 2017). The existence of a past SMBH-IMBH binary in the Galactic center $\gtrsim 10$ million years ago is also supported by the distribution of distances and velocities of hypervelocity stars in the Galactic halo (Cao et al., 2025).
@@
In the case of the Galactic center, current IMBHs could also be detected in less massive, dense stellar associations. Two candidate stellar associations have been identified – IRS 13E (Maillard et al., 2004; Schödel et al., 2005; Tsuboi et al., 2017, 2019; Peißker et al., 2023, 2024) and IRS 1W (Hosseini et al., 2024) at projected distances from Sgr A* of $\sim 0.13\,{\rm pc}$ and $\sim 0.24$ pc, respectively – both of which could host an IMBH of $\sim 10^{4}\,M_{\odot}$ that would make them more stable against tidal disruption by the central supermassive black hole, although some studies tend to disfavor the IMBH presence (Schödel et al., 2005; Fritz et al., 2010; Pavlík et al., 2024) or propose alternative models for the dark mass, such as a dark star cluster of stellar-mass black holes (Banerjee & Kroupa, 2011). These IMBHs, assuming their presence for the purpose of this study, are expected to be in the low-luminosity quiescent state characterized by a radiatively inefficient accretion flow (RIAF) with the spectral energy distribution peaking in the mid-infrared domain (Peißker et al., 2024; Hosseini et al., 2024). A hot RIAF surrounding massive black holes is characteristic of wind-fed accretion provided by host stellar clusters, such as for the case of Sgr A* embedded within the dense nuclear star cluster (Ressler et al., 2018, 2019b). On the other hand, an IMBH could host a dense, optically thick accretion disk if it has recently disrupted a star; such a disk would be revealed by an intense thermal soft X-ray emission with the luminosity reaching a fraction of the Eddington luminosity,
@@
In this paper, we study the scenario of a compact stellar association bound to an IMBH, e.g., a stellar cluster inspiralling towards the Galactic center (Hansen & Milosavljević, 2003). Such a cluster undergoes a continuous tidal disruption as it approaches the SMBH due to dynamical friction and may be representative of IRS 13E, a concentration of a handful of massive early-type stars close to Sgr A* (Peißker et al., 2023, 2024). The constraints on the IMBH mass and IRS 13E distance from the SMBH limit the size of the region that is stable against tidal disruption. This size can be estimated by the tidal (Hill) radius, $r_{\rm H}\simeq d_{\rm IRS13E}(M_{\bullet}/3M_{\rm SgrA*})^{1/3}\sim 17.6\times 10^{-3}\,{\rm pc}(d_{\rm IRS13E}/0.13\,{\rm pc})(M_{\bullet}/3\times 10^{4}\,M_{\odot})^{1/3}(M_{\rm SgrA*}/4\times 10^{6}\,M_{\odot})^{-1/3}$, which is comparable to the projected radii of six early-type stars from the center of the IRS 13E cluster (associated with the E3 source; Peißker et al., 2023). Despite a small number of stars, the region has a very high mean stellar density of $\overline{n}_{\star}\sim 3\times 10^{5}\,{\rm pc^{-3}}$, implying a mean mutual distance among stars of only $\sim 3000$ au, which is expected to result in ongoing wind-wind collisions, leading to shocks and production of hot, X-ray emitting gas. Therefore, such a stellar system provides the motivation for a detailed exploration of the observability of the hypothetical, embedded IMBH and general properties of the cluster inflow-outflow dynamics.
@@
A cluster bound by an IMBH in the central parsec of the galaxy would have a lifetime $\lesssim$ $100{,}000$ years due to a short dynamical friction timescale (Hosseini et al., 2024). An inspiralling cluster is generally unstable and is expected to disrupt with or without the IMBH while interacting with the SMBH and other stars in the NSC (Pavlík et al., 2024). However, the scenario of a compact group of stars surrounding the IMBH is still highly relevant since the IMBH on its way towards the Galactic center can also tidally capture stars from its surroundings (Madau & Rees, 2001), even while it is losing some of them. Furthermore, the scenario not only captures the physics of IRS 13E-like associations in the Galactic center but is also related to the very cores of globular clusters where a handful of stars is expected to orbit hypothetical IMBHs on tight orbits (Häberle et al., 2024).
@@
The mutual interplay of massive stellar winds has been extensively studied in binary systems using hydrodynamical simulations, where high-density outflows create strong shocks and radiative cooling layers (Stevens et al., 1992; Parkin & Pittard, 2008; Pittard, 2009; Parkin et al., 2011; Parkin & Gosset, 2011; Lamberts et al., 2011, 2012; van Marle et al., 2011; Kee et al., 2014; Hendrix et al., 2016; Calderón et al., 2020a). In particular, the wind-wind interface is prone to a thin-shell cooling instability that drives dense clump formation. These clumps can be intermittently stripped or disrupted before reaching the companion star. A special class of binary systems is X-ray binaries where a wind-blowing star provides the material for accretion by a compact companion. This compact companion then focuses the stellar wind into a gaseous tail due to its gravitational field (Hadrava & Čechura, 2012; Čechura & Hadrava, 2015; Čechura et al., 2015).
@@
In between these two regimes, numerical modeling has seldom targeted compact clusters with an embedded IMBH. Compared to binary systems, these clusters feature overlapping wind–wind collision fronts and more complex orbital configurations, yet their sub-pc extent and smaller black-hole masses clearly set them apart from the parsec-scale environment of Sgr A*. Investigating this regime fills a key gap, determining whether wind-fed accretion processes remain similarly inefficient and clumpy and whether the accretion onto a possible IMBH candidate might be detectable above the emission from the colliding stellar winds themselves.
1.1 Structure of the paper
This manuscript is organized as follows. In Sect. 2 we describe the numerical scheme applied to study the HD and MHD of interacting stellar winds around the IMBH, including the radiative cooling function, chemical abundances, and the detailed simulation setup. Subsequently, in Sect. 3, we present the main results of 3D MHD and HD simulations, including plane-parallel cuts of different simulation runs and their associated radial profiles of fluid quantities. Sect 4 follows, where we analyze accretion flow variability and periodicities, and compare the predicted X-ray surface-brightness images with observations of IRS 13E, a candidate stellar system potentially hosting an IMBH close to the Galactic center. We put the presented simulation results in the context of previously published models of wind-fed accretion and discuss IMBH accretion regimes in Sect. 5. We also discuss the observability of IMBHs fed by stellar winds within their host clusters. Finally, we conclude with Sect. 6.

### v2

@@ Abstract
Context. Intermediate-mass black holes (IMBHs) have remained elusive in observations, with only a handful of tentative detections. One promising location for the existence of IMBHs is in dense stellar clusters such as IRS 13E in the Galactic Center. Such systems are thought to be fed by stellar winds from massive Wolf–Rayet (WR) stars. Understanding the accretion dynamics in wind-fed IMBHs in such clusters is important for predicting observational signatures. These systems are, however, relatively unexplored theoretically. Understanding the interplay between IMBHs and the surrounding stars—especially through the high-velocity, dense stellar winds of WR stars—is essential for clarifying IMBH dynamics within such environments and explaining why these objects continue to evade unambiguous detection.
@@
Methods. We conduct three-dimensional magnetohydrodynamic and hydrodynamic simulations in which each WR star represents a source term in mass, momentum, energy, and magnetic field. A metallicity-dependent cooling function accounts for radiative energy losses. By varying the cluster’s geometry, magnetization, and chemical abundance, we characterize the resulting flow structures, accretion rates, and X-ray luminosity corresponding to the cluster and the IMBH.
Results. In all configurations, the accretion rate onto the IMBH is up to five orders of magnitude lower than the total mass-loss rate from the cluster’s WR stars. High-velocity wind–wind collisions generate turbulent, shock-heated outflows that expel most injected gas. While enhanced cooling in high-metallicity runs fosters dense clump formation, these clumps typically do not reach the black hole. The integrated X-ray emission is dominated by colliding stellar winds, rendering the IMBH’s radiative signature elusive. Intermittent, quasi-periodic variations in inflow rates are driven by close stellar passages, leading to short-lived accretion enhancements or flares that nonetheless remain difficult to detect against the dominant wind emission.
Conclusions. Despite continuous mass injection from dense stellar winds, these compact systems exhibit outflow-dominated flows and weak, though strongly variable black-hole accretion, naturally explaining the low detectability of IMBHs in such settings. The observed quasi-periodic or flaring accretion episodes are overwhelmed by the luminous shock-heated winds, making unambiguous observational identification of the IMBH challenging with current X-ray instruments. Nevertheless, our results provide a framework for interpreting current data and developing future observational strategies to unveil IMBHs in dense stellar systems.
@@ 1 Introduction
Intermediate-mass black holes (IMBHs) with masses in the range $M_{\bullet}\sim 10^{2}-10^{5}$ solar masses ($M_{\odot}$) have largely eluded detection, leaving a noticeable gap in the mass distribution of black holes (see Greene et al., 2020, for a review). At the same time, comprehending their formation, evolution, and occurrence across cosmic history is crucial for understanding black-hole growth and galaxy evolution. So far, there have been a few confirmed detections towards the upper IMBH mass range in lower-mass galaxies (e.g., NGC 4395 with $M_{\bullet}\sim 10^{4}-10^{5}\,M_{\odot}$; Filippenko & Ho, 2003; Peterson et al., 2005; Pandey et al., 2024). Toward the lower mass range, there was the first gravitational wave detection of the formation of a $\sim 140\,M_{\odot}$ black hole (GW190521; Abbott et al., 2020), which falls in the pair-instability mass gap (Woosley & Heger, 2021). The most massive of the gravitational-wave merger products has been a $\sim 225\,M_{\odot}$ IMBH (GW231123; The LIGO Scientific Collaboration et al., 2025), which formed the merger of a $\sim 100\,M_{\odot}$ and $\sim 140\,M_{\odot}$ black hole, both high-spinning. This finding supports the hierarchical merger channel for producing black holes of several $\sim 100\,M_{\odot}$.
There are several ways that an IMBH could form, depending on the cosmic time at formation and the surrounding environment, both of which influence its subsequent growth. Three formation channels can be distinguished: (i) primordial/cosmological origin due to the collapse of population III stars (Madau & Rees, 2001) or a direct gas cloud collapse (Loeb & Rasio, 1994; Eisenstein & Loeb, 1995; Begelman et al., 2006; Latif et al., 2013); (ii) consecutive mergers of stellar-mass black holes in dense stellar systems, such as globular clusters (Miller & Hamilton, 2002; Gültekin et al., 2004; Fujii et al., 2024); and (iii) runaway collisions and mergers of massive stars in dense stellar clusters that subsequently collapse into an IMBH (Portegies Zwart & McMillan, 2002). Despite the uncertainties, several theoretical models support (iii), in particular for globular clusters (Miller & Hamilton, 2002; Portegies Zwart & McMillan, 2002; Gültekin et al., 2004; Freitag et al., 2006). Once a sufficiently massive IMBH seed forms ($50\,M_{\odot}$; Miller & Hamilton, 2002), it is expected to be retained within the cluster and undergo only a low-amplitude Brownian-like motion (Chatterjee et al., 2002a, b).
@@
Even granting those assumptions, however, both the occupancy fraction and mass ratio for clusters are uncertain (see e.g. Šubr et al., 2019) due to uncertainties in the initial cluster concentration (Giersz et al., 2015), which influences whether the hypothetical IMBH formation proceeds rapidly or more gradually. Nevertheless, several pieces of observational evidence support that some massive stellar clusters host an IMBH. For instance, the observed off-nuclear X-ray sources and transients in galactic halos with peak X-ray luminosities of $\sim 10^{43}\,{\rm erg\,s^{-1}}$ are consistent with the emission of IMBHs (with masses on the order of $\sim 10^{4}\,M_{\odot}-10^{5}\,M_{\odot}$) powered by the tidal disruption of stars, which supports their association with lower-mass stellar systems like stellar clusters or dwarf galaxies (Lin et al., 2018; Jin et al., 2025; Zhang et al., 2025). There is strong evidence that at least two globular clusters in the Milky Way host IMBHs based on dynamical grounds: $\omega$ Centauri (Häberle et al., 2024) and 47 Tucanae (NGC 104, Kızıltan et al. 2017), with IMBH masses of $\gtrsim 8200\,M_{\odot}$ and $2300_{-850}^{+1500}\,M_{\odot}$, respectively, though several studies point out potential caveats and systematic effects that can influence the central mass constraints (Zocchi et al., 2017, 2019; Bañares-Hernández et al., 2025).
@@
Like the Galactic Center in our own galaxy, Galactic nuclei are especially promising regions for detecting IMBHs. Due to the short dynamical relaxation timescale, as many as $\gtrsim 50(M_{\bullet}/150\,M_{\odot})^{1/2}$ IMBHs could be concentrated towards the supermassive black hole (SMBH, Madau & Rees 2001). These black holes would have been formed primordially or in globular clusters before undergoing a rapid inspiral caused by dynamical friction. The infall rate is expected to decrease with cosmic time as $\propto t^{-1/2}$ and the current rate is estimated to be $\sim 2$ IMBHs per gigayear per galaxy (Madau & Rees, 2001).
@@
In addition, IMBHs up to $\lesssim 10^{4}\,M_{\odot}$ can also form in situ within the dense nuclear star clusters (NSCs) surrounding SMBHs (Neumayer et al., 2020) via black hole–star collisions (Rose et al., 2022; Haas et al., 2025) and black hole-black hole mergers (Fragione et al., 2022). Such IMBHs may be detected indirectly when they perturb the inner accretion flow, causing quasiperiodic accretion and outflow phenomena (Suková et al., 2021; Pasham et al., 2024; Zajaček et al., 2024), through spin-orbit coupling-induced jet precession (Britzen et al., 2023; von Fellenberg et al., 2023; Ressler et al., 2024), or occasionally as X-ray sources when they disrupt a star (Lin et al., 2018; Jin et al., 2025) or encounter a denser molecular cloud (Seepaul et al., 2022; Hu et al., 2024). Furthermore, recent observations suggest that some high-velocity compact clouds (HVCCs) in the Galactic center, such as CO–0.40–0.22, may be influenced by gravitational interactions with inactive IMBHs, highlighting their potential role in the dynamics of molecular clouds (Takekawa et al., 2017). The existence of a past SMBH-IMBH binary in the Galactic center $\gtrsim 10$ million years ago is also supported by the distribution of distances and velocities of hypervelocity stars in the Galactic halo (Cao et al., 2025). The timing of the Sgr A*-IMBH binary also agrees with the merger scenario, which can explain the origin of fast-moving S stars. The SMBH-IMBH merger naturally creates an apse-aligned eccentric disk due to the gravitational-wave recoil kick of the SMBH after the merger. The eccentric disk drives the inner stars into highly eccentric, inclined orbits, consistent with the orbital properties of S stars. According to Akiba et al. (2025) Sgr A* could have merged with the IMBH of $2^{+3}_{-1.2}\times 10^{5}\,M_{\odot}$ within the last 10 million years.
@@
In the case of the Galactic center, current IMBHs could also be detected in less massive, dense stellar associations. Two candidate stellar associations have been identified – IRS 13E (Maillard et al., 2004; Schödel et al., 2005; Tsuboi et al., 2017, 2019; Peißker et al., 2023, 2024) and IRS 1W (Hosseini et al., 2024) at projected distances from Sgr A* of $\sim 0.13\,{\rm pc}$ and $\sim 0.24$ pc, respectively – both of which could host an IMBH of $\sim 10^{4}\,M_{\odot}$ that would make them more stable against tidal disruption by the central supermassive black hole, although some studies tend to disfavor the IMBH presence (Schödel et al., 2005; Fritz et al., 2010; Pavlík et al., 2024) or propose alternative models for the dark mass, such as a dark star cluster of stellar-mass black holes (Banerjee & Kroupa, 2011). These IMBHs, assuming their presence for this study, are expected to be in the low-luminosity quiescent state characterized by a radiatively inefficient accretion flow (RIAF) with the spectral energy distribution peaking in the mid-infrared domain (Peißker et al., 2024; Hosseini et al., 2024). A hot RIAF surrounding massive black holes is characteristic of wind-fed accretion provided by host stellar clusters, such as for the case of Sgr A* embedded within the dense nuclear star cluster (Ressler et al., 2018, 2019b). On the other hand, an IMBH could host a dense, optically thick accretion disk if it has recently disrupted a star; such a disk would be revealed by an intense thermal soft X-ray emission with the luminosity reaching a fraction of the Eddington luminosity,
@@
In this paper, we study the scenario of a compact stellar association bound to an IMBH, e.g., a stellar cluster inspiralling towards the Galactic center (Hansen & Milosavljević, 2003). Such a cluster undergoes a continuous tidal disruption as it approaches the SMBH due to dynamical friction and may be representative of IRS 13E, a concentration of a handful of massive early-type stars close to Sgr A* (Peißker et al., 2023, 2024). The constraints on the IMBH mass and IRS 13E distance from the SMBH limit the size of the stable region against tidal disruption. This size can be estimated by the tidal (Hill) radius, $r_{\rm H}\simeq d_{\rm IRS13E}(M_{\bullet}/3M_{\rm SgrA*})^{1/3}\sim 17.6\times 10^{-3}\,{\rm pc}(d_{\rm IRS13E}/0.13\,{\rm pc})(M_{\bullet}/3\times 10^{4}\,M_{\odot})^{1/3}(M_{\rm SgrA*}/4\times 10^{6}\,M_{\odot})^{-1/3}$, which is comparable to the projected radii of six early-type stars from the center of the IRS 13E cluster (associated with the E3 source; Peißker et al., 2023). Despite a small number of stars, the region has a very high mean stellar density of $\overline{n}_{\star}\sim 3\times 10^{5}\,{\rm pc^{-3}}$, implying a mean mutual distance among stars of only $\sim 3000$ au, which is expected to result in ongoing wind-wind collisions, leading to shocks and production of hot, X-ray emitting gas. Therefore, such a stellar system motivates a detailed exploration of the observability of the hypothetical, embedded IMBH and general properties of the cluster inflow-outflow dynamics.
@@
A cluster bound by an IMBH in the central parsec of the galaxy would have a lifetime $\lesssim$ 100,000100{,}000 years due to a short dynamical friction timescale (Hosseini et al., 2024). An inspiralling cluster is generally unstable and is expected to disrupt with or without the IMBH while interacting with the SMBH and other stars in the NSC (Pavlík et al., 2024). However, the scenario of a compact group of stars surrounding the IMBH is still highly relevant since the IMBH, on its way towards the Galactic center, can also tidally capture stars from its surroundings (Madau & Rees, 2001), even while it is losing some of them. Furthermore, the scenario not only captures the physics of IRS 13E-like associations in the Galactic center but is also related to the very cores of globular clusters where a handful of stars is expected to orbit hypothetical IMBHs on tight orbits (Häberle et al., 2024).
@@
The mutual interplay of massive stellar winds has been extensively studied in binary systems using hydrodynamical simulations, where high-density outflows create strong shocks and radiative cooling layers (Stevens et al., 1992; Parkin & Pittard, 2008; Pittard, 2009; Parkin et al., 2011; Parkin & Gosset, 2011; Lamberts et al., 2011, 2012; van Marle et al., 2011; Kee et al., 2014; Hendrix et al., 2016; Calderón et al., 2020a). In particular, the wind-wind interface is prone to a thin-shell cooling instability that drives dense clump formation. These clumps can be intermittently stripped or disrupted before reaching the companion star. A special class of binary systems is X-ray binaries, where a wind-blowing star provides the material for accretion by a compact companion. This compact companion then focuses the stellar wind into a gaseous tail due to its gravitational field (Hadrava & Čechura, 2012; Čechura & Hadrava, 2015; Čechura et al., 2015).
@@
Numerical modeling seldom targets compact clusters with an embedded IMBH between these two regimes. Compared to binary systems, these clusters feature overlapping wind–wind collision fronts and more complex orbital configurations. Nevertheless, their sub-pc extent and smaller black-hole masses clearly set them apart from the parsec-scale environment of Sgr A*. Investigating this regime fills a key gap, determining whether wind-fed accretion processes remain similarly inefficient and clumpy and whether the accretion onto a possible IMBH candidate might be detectable above the emission from the colliding stellar winds.
This manuscript is organized as follows. In Sect. 2 we describe the numerical scheme applied to study the HD and MHD of interacting stellar winds around the IMBH, including the radiative cooling function, chemical abundances, and the detailed simulation setup. Subsequently, in Sect. 3, we present the main results of 3D MHD and HD simulations, including plane-parallel cuts of different simulation runs and their associated radial profiles of fluid quantities. Sect. 4 follows, where we analyze accretion flow variability and periodicities, and compare the predicted X-ray surface-brightness images with observations of IRS 13E, a candidate stellar system potentially hosting an IMBH close to the Galactic center. We put the presented simulation results in the context of previously published models of wind-fed accretion and discuss IMBH accretion regimes in Sect. 5. We also discuss the observability of IMBHs fed by stellar winds within their host clusters. Finally, we conclude with Sect. 6.